Assume: Demand Curve: QD = 80 – 10P; and Supply Curve: QS = 10P 7. Given the information derived above, identify on the graph consumer surplus and producer surplus for each situation as well as deadweight loss, if any. INCLUDE A GRPAH c. Government imposes a maximum price of $2.00. INCLUDE A GRAPH
Answer #1

Government imposes a maximum price of $2

New Equilibrium price= $2

New Equilibrium Quantity= 20

Consumer surplus=(0.5*20*(8-6))+(6-2)*20=20+80= $100

Producer surplus= 0.5*2*20= $20

Deadweight loss= 0.5(6-2)*(40-20)= $40

Price GON 006 TIF Supply 5 Consumeris Surplus Deadweight Loss Producer Surplus Price Ceiling Demand 10 20 30 40 50 60 70 80 -